# Interpretable Measurement of CNN Deep Feature Density using Copula and the Generalized Characteristic Function

### v2

@@
Interpretable Measurement of CNN Deep Feature Density using Copula and the Generalized Characteristic Function
@@
{adjustwidth}
25pt25pt We present a novel empirical approach toward measuring the Probability Density Function (PDF) of the deep features of Convolutional Neural Networks (CNNs). Measurement of the deep feature PDF is a valuable problem for several reasons. Notably, a. Understanding the deep feature PDF yields new insight into deep representations. b. Feature density methods are important for tasks such as anomaly detection which can improve the robustness of deep learning models in the wild. Interpretable measurement of the deep feature PDF is challenging due to the Curse of Dimensionality (CoD), and the Spatial intuition Limitation. Our novel measurement technique combines copula analysis with the Method of Orthogonal Moments (MOM), in order to directly measure the Generalized Characteristic Function (GCF) of the multivariate deep feature PDF. We find that, surprisingly, the one-dimensional marginals of non-negative deep CNN features after major blocks are not well approximated by a Gaussian distribution, and that these features increasingly approximate an exponential distribution with increasing network depth. Furthermore, we observe that deep features become increasingly independent with increasing network depth within their typical ranges. However, we surprisingly also observe that many deep features exhibit strong dependence (either correlation or anti-correlation) with other extremely strong detections, even if these features are independent within typical ranges. We elaborate on these findings in our discussion, where we propose a new hypothesis that exponentially infrequent large valued features correspond to strong computer vision detections of semantic targets, which would imply that these large-valued features are not outliers but rather an important detection signal.
@@ 1 Introduction
Convolutional Neural Networks (CNN) have revolutionized the performance of image analysis tasks including image classification, semantic segmentation, object detection, and image and video synthesis [34, 7, 9, 32]. At the time of writing, CNNs continue to play a domenent role as image feature encoders for state-of-the-art techniques including several prominent Vision-Language Models (VLMs) [22, 25, 20] as well as diffusion models for image generation [28, 33]. Nevertheless, the extraordinary complexity of CNNs has coined the nick-name of black-box, that learns an uninterpretable and high-dimensional feature representation. This work contributes toward improving our understanding of the learned representation by measuring the statistical characteristics of deep CNN features through high-dimensional statistical analysis techniques including copula analysis, and the Method of Orthogonal Moments (MOM) to obtain the Generalized Characteristic Function (GCF). Our approach is unique because it provides a novel and interpretable probability density estimate of CNN features, without making any rigid parametric assumptions that may be unjustified, and without altering the CNN feature representation which may reduce accuracy. Furthermore, as an empirical technique, we can gain greater insight by plotting and analyzing the marginal and interdependence components of the feature copula density PDF which can lead to improved understanding of the statistical behavior of the native CNN feature space. We want to measure and observe the probability density of deep features for popular CNN architectures, as we believe that this will lay the groundwork for future feature-density analysis methods that can identify stronger parametric assumptions, thereby leading to improved understanding of the feature space as well as improved probability density estimation of the distribution of deep CNN features.
Measuring the probability density of deep CNN features has many practical applications including Out-of-Distribution detection [19, 21, 36, 13, 18], adversarial detection [19, 18], domain generalization [3], federated learning [31], and semi-supervised learning [23]. It is anticipated that if one knows the density distribution of deep features, that it would be possible to use outlier detection to statistically distinguish legitimate inlier data from anomalous outlier data, thereby providing a greater level of model robustness to data that is out of model scope [8, 19, 21]. Furthermore, methods that can model both the conditional as well as the joint distribution meet the technical definition of generative methods rather than purely discrimination methods [14, 2, 15, 26]. Generative methods have many theoretical advantages because they enable complex inferencing tasks regarding the distribution of unlabeled samples in addition to performing basic discrimination tasks.
Several works have attempted to construct deep generative architectures with explicit feature representations that enable exact or approximate PDF estimates [14, 2, 15, 26]. Such techniques include variational auto-encoders [14], as well as autoregressive and normalizing flows [2, 15, 26]. Nevertheless, open questions remain regarding the applicability of deep generative PDFs for anomaly detection for several reasons including the often unaccounted for effects of reparameterization [18] and class-imbalance [13], as well as questions regarding the empirical validity of the underlying ’typical set’ hypothesis [35] stating that there is a discernible difference between the statistical properties of typical valued in-distribution and out-of-distribution features [19, 21, 36]. We believe that further empirical analysis, including alternate ways of obtaining density estimates will be valuable toward studying these observed phenomena.
@@
We take a purely empirical approach toward accurate estimation of deep CNN feature density. An important contribution of our approach is that we do so without unverified parametric or linearity assumptions, and without altering the CNN architecture or feature representation which may reduce accuracy. Several prior works have assumed the CNN features follow a Multi-Variate Gaussian (MVG) [23, 36, 19, 27, 36]. But the motivation for MVG traces back to a handful of histogram plots of the penultimate critic features in the supplemental materials of Lee et al. [19]. There has never been a verification that MVG is suitable to the deeper intermediate features (non-penultimate), which is especially important because in most architectures, deep features undergo ReLU activation which deactivates (zeros out) any negative valued features, thereby enforcing non-negativity. Max-pooling moreover retains the strongest positive features, throwing away weaker less-positive detections. Thus it is important to look at and measure the distribution of non-negative features as these are the only features that have an impact on subsequent activations. We observe that these non-negative deep features are highly non-Gaussian, and increasingly approximate an exponential distribution with increasing network depth. We furthermore observe that at adequate network depth, many deep features exhibit non-linear dependence, with typical feature values showing statistical independence for typical (non-extreme) values, but strong statistical dependence of extreme values. In other words, we observe that two deep features may be uncorrelated within their typical ranges, but if one feature is extremely large or small, then other features may show strong correlation or anti-correlation in their extreme values.
This work also contributes to an active and growing body of ongoing empirical work toward improved CNN feature understanding [1, 3, 6, 11, 24, 29]. Several recent works have attempted to visualize CNN features, in order to determine under what circumstances a model may learn or retain features that are spurious (correlated) versus invariant (causal) to semantic meaning [3, 24, 29]. We do not distinguish spurious versus invariant features in this work, however we do observe strong correspondence between extreme-valued features, suggesting that many deep features within intermediate representations may be learning correlated (or anti-correlated) semantic concepts with a statistical backdrop of uncorrelated detection noise. Allen-Zhu and Li [1] propose a multi-view hypothesis that intermediate deep CNN features may learn to detect specific views of target objects. Based on our results, we propose a follow-on hypothesis which we coin the exponential view hypothesis. That is, at adequate network depth, the distribution of feature detections for any given semantic target view follows exponential distribution. That is, prevalent small-value features correspond to a lack of detection of their respective target view, whereas rare large-value features correspond to a strong detection of the target view. As the actual target often corresponds to only a small portion of the image frame, the strong-valued detections are statistically rare occurrences but represent important computer vision signal, whereas the lack of detection, although more statistically prevalent, is primarily a noise signal. Our exponential view hypothesis offers a new perspective on the possible empirical shortcomings of the typical set hypothesis. It logically follows that the typical set may be accidentally eliminating the most important computer vision signal by excluding the extreme-value features from consideration in the typical set. We discuss this hypothesis further in the discussion section.

### v3

# Interpretable Estimation of CNN Deep Feature Density using Copula and the Generalized Characteristic Function

@@
Interpretable Estimation of CNN Deep Feature Density using Copula and the Generalized Characteristic Function
@@
We present a novel empirical approach toward estimating the Probability Density Function (PDF) of the deep features of Convolutional Neural Networks (CNNs). Estimating the PDF of deep CNN features is an important task, because it will yield new insight into deep representations. Moreover, characterizing the statistical behavior has implications for the feasibility of promising downstream tasks such as density based anomaly detection. Expressive, yet interpretable estimation of the deep feature PDF is challenging due to the Curse of Dimensionality (CoD) as well as our limited ability to comprehend high-dimensional inter-dependencies. Our novel estimation technique combines copula analysis with the Method of Orthogonal Moments (MOM), in order to directly estimate the Generalized Characteristic Function (GCF) of the multivariate deep feature PDF. We find that the one-dimensional marginals of non-negative deep CNN features after major blocks are not well approximated by a Gaussian distribution, and that the features of deep layers are much better approximated by the Exponential, Gamma, and/or Weibull distributions. Furthermore, we observe that deep features become increasingly long-tailed with network depth, although surprisingly the rate of this increase is much slower than theoretical estimates. Finally, we observe that many deep features exhibit strong dependence (either correlation or anti-correlation) with other extremely strong detections, even if these features are independent within typical ranges. We elaborate on these findings in our discussion, where we hypothesize that the long-tail of large valued features corresponds to the strongest computer vision detections of semantic targets, which would imply that these large-valued features are not outliers but rather an important detection signal.
@@ 1 Introduction
Convolutional Neural Networks (CNN) have revolutionized the performance of image analysis tasks including image classification, semantic segmentation, object detection, and image and video synthesis yuan2016feature ; goodfellow2020generative ; hao2020brief ; xing2023survey . At the time of writing, CNNs continue to play a dominant role as image feature encoders for state-of-the-art techniques including several prominent Vision-Language Models (VLMs) long2024hybrid ; radford2021learning ; li2021align as well as diffusion models for image generation rombach2022high ; yang2023diffusion . Nevertheless, the extraordinary complexity of CNNs has coined the nick-name of black-box, that learns an uninterpretable and high-dimensional feature representation. Although there has been progress toward understanding the semantic meaning of features, there has been much less progress toward understanding the high-dimensional probability distribution that underlies the statistical behavior of deep CNN features zhang2025systematic ; vladimirova2019bayesian ; feldman2020short ; feldman2020discovering ; allen2023towards ; chen2024understanding ; giraldo2019integrating ; hermann2020shapes ; qiu2024complexity ; shwartz2017opening .
@@
What is the distribution of deep CNN features? The goal of this work is to improve our understanding of the statistical behavior of the learned representation of deep CNN features through high-dimensional statistical analysis techniques including copula analysis, and the Method of Orthogonal Moments (MOM) sklar1959copula ; nelsen2006copula ; yudell1975mathematical . Rather than assuming a parametric distribution apriori, our approach is to directly estimate the Generalized Characteristic Function (GCF) hansen1982large of the copula interdependence term nelsen2006copula . Our copula+GCF approach provides a novel and interpretable probability density estimate of CNN features, without rigid parametric assumptions that may be unjustified, and without altering the CNN feature representation which may reduce accuracy. As an empirical technique, we can gain greater insight by plotting and analyzing the marginal and interdependence components of the feature copula density PDF. We want to estimate and observe the probability density of deep features for popular CNN architectures, as we believe that this will lay the groundwork for future feature-density analysis methods that can identify stronger parametric assumptions, thereby leading to improved understanding of the feature space and its statistical behavior.
@@
Several prior works have assumed the CNN features follow a Multi-Variate Gaussian (MVG) majurski2024method ; zhu2022boosting ; lee2018simple ; rippel2021modeling ; zhu2022boosting . But the motivation for MVG traces back to a handful of histogram plots of the penultimate critic features in the supplemental materials of Lee et al. lee2018simple . There has never been a verification that MVG is suitable to the deeper intermediate features (non-penultimate), which is especially important because in most architectures, deep features undergo ReLU activation which deactivates (zeros out) any negative valued features, thereby enforcing non-negativity. Max-pooling moreover retains the strongest positive features, throwing away weaker less-positive detections. Thus it is important to look at and measure the distribution of non-negative features as these are the only features that have an impact on subsequent activations. We observe that these non-negative deep features are highly non-Gaussian, and increasingly approximate long-tailed distributions with increasing network depth. We furthermore show that at adequate network depth, many deep features exhibit non-linear dependence, with typical feature values showing statistical independence for typical (non-extreme) values, but strong statistical dependence of larger extreme values, a phenomenon known as upper tail dependence nelsen2006copula .
Several works have attempted to construct deep generative architectures with explicit feature representations that enable exact or approximate PDF estimates kingma2013auto ; caterini2021variational ; kingma2016improved ; rezende2015variational . Such techniques include Variational Auto-Encoders (VAE) kingma2013auto , as well as autoregressive and normalizing flows caterini2021variational ; kingma2016improved ; rezende2015variational . VAE provides a very coarse estimate of the PDF of a deep embedding layer through the incorporation of approximate KL-divergence constraints with an assumed parametric prior (typically standard Gaussian) kingma2013auto . Unlike VAE, our approach provides a fine grained estimate of the PDF for multiple feature layers simultaneously and without rigid parametric constraints. Normalizing flows take a very different approach of propagating feature probability density layer-by-layer over a series of reversible layers by means of the change of variables formula kingma2016improved ; rezende2015variational . Like our technique, normalizing flows provide a fine grained PDF for multiple layers simultaneously, however, our approach has two advantages: 1. The need for reversible layers means that normalizing flows cannot be applied to popular CNN architectures such as ResNet and VGG without substantial redesign. 2. Our approach is interpretable in the sense that we can make visually understandable plots describing the shape of the feature marginal and interdependence terms. As such, our goal is not only to model the probability density and achieve a good fit, but moreover to represent the probabilities in a way that can facilitate visual exploration of the statistical behavior.
It has been proposed that estimating the probability density of deep CNN features could have many practical downstream applications including Out-of-Distribution (OOD) detection lee2018simple ; liu2020energy ; zhu2022boosting ; jiang2023detecting ; le2021perfect , adversarial detection lee2018simple ; le2021perfect , domain generalization chen2024understanding , and federated learning sun2023feature . It has been anticipated that if one knows the density distribution of deep features, that it would be possible to use outlier detection to statistically distinguish legitimate inlier data from anomalous outlier data, thereby providing a greater level of model robustness to data that is out of model scope guerin2023out ; lee2018simple ; liu2020energy . Perhaps the most straightforward of these downstream tasks is anomaly detection (including OOD detection), which was first proposed by Bishop bishop1993novelty as an application of neural network feature density. Density based anomaly detection (outside of deep learning) is a traditional approach with historical success chandola2009anomaly . Recently, density based anomaly detectors have been implemented and report good performance on evaluation tasks cook2024feature ; liu2022unsupervised .
But other recent papers cast doubt on whether density based anomaly detection is even possible using deep CNN features nalisnick2018deep ; lee2018simple ; liu2020energy ; zhu2022boosting ; le2021perfect ; jiang2023detecting ; zhang2021understanding . Nalisnick et al. nalisnick2018deep was the first to observe the strange result that feature density methods, including VAE and normalizing flows, often attribute greater density to the out-of-distribution samples relative to the in-distribution samples. The research community is still searching for an answer as to why traditional density based anomaly detection methods have fallen short when applied to deep features, and several competing hypotheses have been proposed. Hypotheses include the often unaccounted for effects of reparameterization le2021perfect and class-imbalance jiang2023detecting , as well as questions regarding the empirical validity of the underlying ’typical set’ hypothesis zhang2021understanding stating that there is a discernible difference between the statistical properties of typical valued in-distribution and out-of-distribution features lee2018simple ; liu2020energy ; zhu2022boosting . Although the true reason for this behavior may be difficult to pinpoint, we believe that understanding the statistical behavior of the distribution of deep features plays a vital role in answering this question. If deep CNN features followed the same distribution as traditional anomaly detection datasets, one would expect traditional density-based outlier detection methods to perform well just as they have in the past bishop1993novelty ; chandola2009anomaly . Empirical work to visually interpret and understand the distribution of deep features will provide new insight toward answering this important question. Based on our results, we propose a new Long-tailed view hypothesis, that many deep features exhibit long-tailed marginals for which outliers are likely to occur, and moreover that these outliers correspond to strong computer vision detections of a given target view. Thus, attempting to remove outliers for the purpose of anomaly detection may in fact mistakenly remove the most important detection signal.
There are several reasons to suspect that the marginal distribution of deep CNN features might exhibit a long-tailed distribution, although this phenomenon has not been previously verified through direct analysis of the feature marginals zhang2025systematic ; vladimirova2019bayesian ; feldman2020short ; feldman2020discovering . Long-tailed theory suggests that the real-world image set exhibits many rare instances, be it classification subcategories, or image statistics. Early works have established that the frequency of object sub-categories naturally follows a long tail salakhutdinov2011learning ; zhu2014longtail . It is now understood that a hidden long tail of rare instances persists even within class balanced datasets; i.e. imbalanced subcategories exist even if they are not explicitly labeled as such feldman2020short . Could it be that deep CNNs learn long-tailed deep feature marginals in order to better represent these long-tailed image statistics? At present, there is indirect empirical supporting evidence for this hypothesis. Feldman and Zhang feldman2020discovering provide rigorous empirical evidence that popular CNNs are memorizing rare instances, in order to more accurately model long-tailed image statistics hidden within popular image datasets. Feldman and Zhang feldman2020discovering use a cross-validation approach to identify the long tail of rare instances within the dataset. The CNN performance significantly drops if these rare instances are excluded from training. Moreover, simply fine-tuning the last layer with these rare instances is not adequate to restore model performance. Therefore it is concluded that the memorization of the long-tail of rare instances occurs within the deeper convolution layers. But we wonder: Could the deeper convolution layers exhibit long-tailed feature marginals in order to better represent the long-tail of rare instances? Our empirical analysis of marginals shows that for all of the models and datasets under consideration, that at least one such deeper layer exhibits long-tailed feature marginals.
The strongest theoretical evidence in support of long-tailed deep CNN feature marginals comes from the analysis of Vladimirova et al. vladimirova2019bayesian which proves that Bayesian neural networks (including Bayesian CNNs), exhibit long tailed deep features for Layer 3 and deeper under the assumption of independent unit Gaussian inputs and weight priors. More-specifically, these deep features are sub-Weibull with optimal Weibull tail parameter $\theta=\frac{1}{2}i$ where $i$ is the depth of the CNN layer. Optimal tail parameter $\theta$ is considered to be the largest value of $\theta$ for which the marginal is sub-Weibull. This means conv Layer 1 is sub-Gaussian ($\theta=\frac{1}{2}$), Layer 2 is sub-exponential ($\theta=1$), and subsequent layers are long-tailed (Layer 3: $\theta=1\frac{1}{2}$, Layer 4: $\theta=2$, Layer 5: $\theta=2\frac{1}{2}$ etc.). This theoretical result very clearly shows how long-tailed deep feature marginals can arise from a neural network even in the absence of long-tailed image statistics, as the assumed Gaussian inputs are certainly not long-tailed. Nevertheless, the proof requires strong assumptions on the distribution of weights and inputs, it is not known to what extent this proof may be useful in explaining the empirical distribution of deep feature marginals for popular CNN architectures and image datasets. In our section on Analysis of marginals, we numerically estimate the optimal Weibull $\theta$ parameters for ResNet and VGG architectures in several image datasets, and show that the empirical behavior of $\theta$ is largely inconsistent with the predicted results of Vladimirova et al. vladimirova2019bayesian , although there are some similarities. Notably, the interior layers of the network do indeed show some steady increase in $\theta$ parameter with layer depth consistently across all models and datasets, although the rate of increase is significantly slower than the theoretical $\theta=\frac{1}{2}i$. Moreover, we find that for all models and architectures, at least one sufficiently deep intermediate layer is indeed long-tailed $\theta>1$, but depending on the model and architecture, many layers may also be sub-exponential $\theta<1$. Finally, the optimal $\theta$ value for the deepest conv layer appears to depend heavily on the dataset, with more complex image datasets (CIFAR-100, Imagenette2) exhibiting larger $\theta$ values than simpler image datasets (CIFAR-10, MNIST) suggesting that image statistics also play a role in whether the deepest layers are long-tailed.
A contribution of our work is that we are the first to apply copula analysis toward understanding the statistical behavior of deep CNN features. Moreover, we believe that we are the first to apply copula analysis toward answering any basic research question in the area of deep machine learning. Copula analysis is a popular method for understanding multivariate joint probabilities in domains such as economics kim2020review , physical sciences bhatti2019recent , and healthcare demongeot2013archimedean . Bhatti and Do bhatti2019recent refer to copulas as ”the most powerful tools that can model dependent structures between various complex correlated variables.” Indeed, copula have the ability to not only model the joint density, but also yield visually expressive plots to facilitate greater understanding of the statistical behavior. Nevertheless, the vast majority of works that have attempted to combine copula with deep learning have been highly applied toward the same traditional problem domains for which copula analysis was historically popular huang2022spatio ; ouyang2019modeling ; bedoui2023portfolio ; silva2020copula . The method of Ling et al. ling2020deep is notable as it provides a deep learning approach toward estimating the parameters of an Archimedean copula. However, even the benchmarks within Ling et al. ling2020deep are highly applied toward the traditional financial modeling and econometrics domains. We are unaware of any prior work that applies copulas toward better understanding of the behavior of deep learning models. Our primary goal is to gain empirical insight into the question: what is the probability distribution of deep CNN features?.
Another contribution of our approach is that our novel copula model exhibits greatly improved expressive power because we are the first to compute a GCF for the purpose of representing the copula density nelsen2006copula ; yudell1975mathematical ; hansen1982large . This is a large improvement over standard Archimedean copulas which are inflexible, because the joint copula density depends on only on a single rank correlation coefficient, typically either Spearman’s $\rho$ or Kendall’s $\tau$ nelsen2006copula . For Archimedean copulas, one must choose from a parametric generating function such as the Gumbel, Frank, or Clayton copula generators. The choice of the parametric generator gives some blunt flexibility over the presence of upper and lower tail dependence in the bivariate copula nelsen2006copula . This makes Archimedean copulas quick to compute and easy to interpret, but the rigid parametric assumptions raise the risk of assuming an unjustified joint distribution, especially if one does not perform extensive exploratory analysis beforehand. By comparison, our GCF copula also includes Spearman’s $\rho$, as this fully defines to the orthogonal moment $\mu_{1,1}$ corresponding of the bivariate linear Legendre polynomial term. But in addition to Spearman $\rho$, we also take into account a much larger number of rank-shape statistics as part of the GCF, as indeed all of the moments constitute different rank-shape statistics. This allows us to non-parametrically model the copula density rather than assuming a simple parametric form. Moreover, the GCF easily generalizes to multivariate joint densities, whereas multivariate Archimedean copulas are typically implemented using either a series of bivariate joint copulas, or by assuming only a single rank-order statistic for the entire multivariate copula nelsen2006copula . As such, our proposed copula is far more expressive than the standard Archimedean copula design, and is therefore able to more accurately model the joint probability density, while maintaining the ease of visualization, and incorporating the familiar Spearman’s $\rho$ term within the copula density calculation.
In recent years, orthogonal functions and the method of moments have found a variety of uses in deep feature spaces. These techniques can allow for compression of the feature space, as well as have led to the development of improved loss functions for GANs and domain adaptation. PCA decomposition of deep feature covariane is a viable approach toward model compression while maintaining predictive power garg2019low , orthogonal gradient updates can also preserve prior knowledge and prevent catastrophic forgetting saha2021gradient ; farajtabar2020orthogonal . PCA is however a purely linear method. The method of moments has also found utility as a way to introduce non-linearity within orthogonal deep feature decompositions majurski2024method ; li2015moment ; li2017mmdgan ; Ansari2020cvpr . Notably, the Maximum Mean Discrepancy (MMD) statistic is widely used as a loss function for deep generative models. MMD is in fact mathematically derived from the method of moments by using the kernel trick with Radial Basis Functions (RBF) li2015moment ; li2017mmdgan . Ansari et al. Ansari2020cvpr introduce a Characteristic Function Distance (CFD) based on the Fourier moment series as a fast alternative to MMD loss for generative models. Majurski et al. majurski2024method introduce a method of moments loss function based on the power series which is used as embedding constraint for contrastive learning. Although the method of moments has found utility within deep representation learning, it is almost always used only as a loss function. In theory, the method of moments is also capable of recovering the full probability density of the deep features, but has rarely been applied for this purpose outside of niche applications.
Characteristic functions are in fact a powerful and expressive way of estimating probability density, although this approach has found limited and niche applications hansen1982large ; yu2023characteristic ; choi2020probabilistic ; nolan2013elliptically . The Empirical Characteristic Function (ECF) was first introduced by Feuerverger and Mureika feuerverger1977empirical who advocated its use for many problems. McCulloch mcculloch199613 Argues that financial modeling and time series forecasting would greatly benefit from the ECF estimation. Yu yu2004empirical shows that ECF outperforms Gaussian Mixture Models (GMM) in modeling financial timeseries forecasts. Nolan nolan2013elliptically further analyzes the properties of the ECF and provides numerical routines for estimation of densities as applied to timeseries data. Yu et al. yu2023characteristic combines ECF with probabilistic circuits choi2020probabilistic to construct a powerful hybrid density estimation technique as estimated using UCI benchmark datasets, and found this approach to be competitive with state of the art density estimation techniques. Although ECF are capable of density estimation, their application has been slow to gain traction outside of the financial modeling, and timeseries forecasting community. Moreover, we are unaware of prior works that have incorporated ECF as a means to compute copula density as we have within our methodology.
Our copula+GCF approach also exhibits improved expressive power over ECF in modeling distributions with long-tailed and/or discontinuous marginal densities. The traditional ECF attempts to approximate a probability distribution by estimating its Fourier transform. But the Fourier transform, which is based on sinsin and coscos terms may require an impractical number of frequencies in order to approximate functions with power law asymptotes (as in long-tailed marginals) and/or discontinuities as are prevalent in non-negative distributions. There are also open questions about how best to sample from the ECF. By separating the marginal and interdependence terms, our copula+GCF strategy allows for empirical characteristic approximation of multivariate distributions that would be impractical to sample from using a standard ECF approach. Furthermore, our GCF construction provides a straightforward sampling strategy as it is based on a discreet well ordered series of orthogonal functions, rather than a continuous set of orthogonal functions as in the ECF. As such, we find that our copula+GCF approach performs favorably to the ECF when used to estimate the density of deep CNN features.
Our method is designed to both expressively model the probability density of deep features while also making it easier for a human to visually interpret the statistical behavior of features through copula analysis and visualization of the marginal and interdependence terms. As such, by interpretability we are referring to the ability to construct the probability density function in a way that leads to interpretation of the sample statistics. This is distinct from the more common notion of interpretability in machine learning, including network architectures with the ability to highlight the semantic meaning of features, and/or provide visual explanations of predictions noh2015learning ; kolouri2017explaining ; zhang2018interpretable ; allen2023towards ; chen2024understanding ; hermann2020shapes . High dimensional statistics are notoriously difficult to comprehend by a human, and copula analysis can help with this interpretation. As such, our empirical method provides an expressive ability to reconstruct the probability density of deep features, while also providing a new visual aid for researchers who may be interested in understanding the statistical distribution of feature marginals and interdependence.